Circular No. 5956

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1994F, 1994G, 1994H S. Perlmutter, C. Pennypacker, G. Goldhaber, A. Goobar, R. Pain, B.

Grossan, A. Kim, M. Kim, and I. Small, Lawrence Berkeley Laboratory and the Center for Particle Astrophysics, Berkeley, report three discoveries from a search for pre-maximum-light, high-redshift supernovae by themselves and R. McMahon, Institute of Astronomy, Cambridge; P. Bunclark, D. Carter, and M. Irwin, Royal Greenwich Observatory; M. Postman and W. Oegerle, Space Telescope Science Institute; T. Lauer, National Optical Astronomy Observatory; and J. Hoessel, University of Wisconsin. Following are given the designation, date of first detection, discovery magnitude and telescope (INT = 2.5-m Isaac Newton Telescope; KPNO = 4-m Kitt Peak telescope), supernova position for equinox 1950.0, offsets from the host galaxy’s center, and date of the previous image of the galaxy not showing the supernova (to limiting mag ∼ 24): SN 1994F, Jan. 9, R = 22.0, INT, α = 11^{h}47^{m}25^{s}.15, δ = +10^{o}59′38′′.8, 1′′.1 west, 0′′.2 north, 1993 Dec. 22; SN 1994G, Feb. 13, I = 21.8, KPNO, α = 10^{h}16^{m}17^{s}.38, δ = +51^{o}07′23′′.5, 1′′.4 east, 0′′.1 north, 1994 Jan. 16; SN 1994H, Jan. 8, R = 21.9, INT, α = 2^{h}37^{m}32^{s}.22, δ = −1^{o}46′57′′.5, 1′′.2 west, 0′′.1 south, 1993 Dec. 20. On Jan. 18, spectra of SN 1994F were obtained by J. B. Oke with the Keck Telescope Low Resolution Imaging Spectrograph; the host galaxy redshift is 0.354, and the spectrum of SN 1994F matched that of a type-Ia supernova a week past maximum light. On Mar. 9 and 10, spectra of SN 1994G were obtained by A. Riess, P. Challis, and R. Kirshner at the Multiple Mirror Telescope, in which emission lines of [O II] and [O III] from the host galaxy give a redshift of z = 0.425; the spectrum of the SN 1994G, though noisy, is consistent with a type-I supernova about a week past maximum light. SN 1994H was observed on numerous nights from Jan. 10 to Feb. 16 at the INT, at Kitt Peak by G. Jacoby and others, at the European Southern Observatory by M. Turrato, and at Siding Spring Observatory by M. Dopita; the resulting photometry is consistent with a type-Ia supernova at an implied redshift of ∼ 0.32 (the host galaxy is on the periphery of a cluster with that redshift), with maximum light around Jan. 12.

COMET McNAUGHT-RUSSELL (1993v) Total visual magnitude estimates by R. Keen, Mt. Thorodin, CO (7×50

binoculars): Mar. 10.09 UT, 7.6 (20′ coma); 22.12, 6.6.

1994 March 24 Daniel W. E. Green